QUESTION 3 Calculate the pH of a solution of HCl with a concentration of 3.00 x 10->M. (Significant figures)
Answer #1

pH of a solution is defined as the negative logarthim of hydronium concentration.

pH = -log(H3O+)

The dissociation of HCl in water can be written as

                 HCl + H2O --------> H3O+ + OH-

Since HCl is a strong acid, the concentration of hydronium ion isequals to the concentration of the acid.

pH= -log(3.00*10^-3) = 2.523